# On $t$-structures adjacent and orthogonal to weight structures

### v1

@@ 1.2 Some basics on weight structures
    The class $\underline{C}_{[m,l]}$ is the extension-closure of $\cup_{m\leq j\leq l}\underline{C}_{w=j}$.
    Consequently, $\underline{C}_{w=0}$ is silting if $w$ is bounded.
  9. 9.
@@
Assume that $\mathcal{B},\mathcal{A}$ are strictly full triangulated subcategories of $\underline{C}$.
Then the couple $D=(\mathcal{A},\mathcal{B})$ is a (length $2$) semi-orthogonal decomposition of $\underline{C}$ (or just gives a decomposition of $\underline{C}$) if $\operatorname{Obj}\mathcal{B}\perp\operatorname{Obj}\mathcal{A}$ and $\operatorname{Obj}\mathcal{B}\star\operatorname{Obj}\mathcal{A}=\operatorname{Obj}\underline{C}$.
@@
    The full subcategories of $\underline{C}$ corresponding to $C_{2}$ and $C_{1}$ give a decomposition of $\underline{C}$.
@@
    $(C_{1},C_{2})$ is a weight structure $w$ and ${\underline{Hw}}=\{0\}$.
@@
    $(C_{1},C_{2})$ is a weight structure $w$ and the triangle (1.2.1) is functorially determined by $M$.
@@
    $(C_{1},C_{2})$ is a weight structure $w$ and the triangle (1.2.1) is determined by $M$ up to a non-canonical isomorphism.
@@
The equivalence of conditions 3, 2, and 2 can be easily obtained by means of combining Propositions 3.4(4) and 3.2(1,2) of [BoV19] (pay attention to Remark1.1.6(2) above!).
The equivalence of conditions 1 and 3 is very well-known; see Proposition 2.5 of [Bon24].
@@
Lastly, assume that condition 5 is fulfilled. Then any weight decomposition triangle is clearly a $1$-weight decomposition triangle as well. Consequently, the uniqueness provided by Proposition 1.2.4(4) yields condition 7; cf. Remark 1.2.6 of [Bon22b] where compositions of the corresponding morphisms between decomposition triangles is discussed. ∎
@@
1. Roughly, semi-orthogonal decompositions are the weight structures that are also $t$-structures. One may also characterize them both as shift-stable weight structures and as shift-stable $t$-structures.
2. Note also that this notion is obviously self-dual (that is, permuting the classes in the couple yields a semi-orthogonal decomposition in the category $\underline{C}^{op}$ cf. Proposition 1.2.4(1).
3. Below we will also need the following simple criterion: if $\underline{C}$ is smashing (cf. Theorem 0.1) then $w$ is smashing if and only if the right adjoint functor provided by condition 3 of Proposition 1.2.6 respects coproducts; see Proposition 3.4(5) of [BoV19].
@@ 1.3 On orthogonal and adjacent structures
1. We will say that $w$ and $t$ are orthogonal if $\underline{C}_{w\leq 0}\perp_{\underline{D}}\underline{C}^{\prime}_{t\geq 1}$ and $\underline{C}_{w\geq 0}\perp_{\underline{D}}\underline{C}^{\prime}_{t\leq-1}$.
@@
Let us now relate the latter definition to the notion of adjacent structures introduced in [Bon10a].
@@
  1. 1.
    In Definition 2.5.1(3) of [Bon10b] and §5.2 of [Bon19] orthogonality was defined with respect to an arbitrary duality $\Phi$ between $\underline{C}$ and $\underline{C}^{\prime}$; see Definition 1.1.1. Moreover, in the current paper we (essentially) restrict this definition to dualities provided by Lemma 1.1.2.
  2. 2.
    Proposition 1.3.3 says that our definition of adjacent "structures" is essentially equivalent to the original Definition 4.4.1 of [Bon10a].
@@
Assume that $w$ and $t$ satisfy the assumptions of Definition 1.3.1, $t$ is strictly orthogonal to $w$ (note that this is condition follows from orthogonality whenever $\underline{C}^{\prime}=\underline{C}$; see Theorem 2.3.3(I) below), and the couple $(\underline{C}_{w\leq 0},\underline{C}_{w\geq 0})$ is a semi-orthogonal decomposition (see Definition 1.2.5). Then $(\underline{C}^{\prime}_{t\geq 0},\underline{C}^{\prime}_{t\leq 0})$ is a semi-orthogonal decomposition as well.
@@
Applying Proposition 1.2.6 (see conditions 4 and 1 in it) along with the definition of strict orthogonality we obtain that $\underline{C}^{\prime}_{t\geq 0}=\underline{C}^{\prime}_{t\geq 0}[1]$. Hence $(\underline{C}^{\prime}_{t\geq 0},\underline{C}^{\prime}_{t\leq 0})$ is a semi-orthogonal decomposition indeed; see condition 2 in Proposition 1.2.6. ∎
@@ 2 On virtual $t$-truncations and their relation to orthogonal $t$-structures
In §2.3 we prove that if $\underline{C}^{\prime}\subset\underline{C}$ then for a weight structure $w$ on $\underline{C}$ there exists at most one $t$-structure on $\underline{C}^{\prime}$ (strictly) orthogonal to it; see Theorem 2.3.3. Moreover, if ${\underline{C}^{0}}\subset\underline{C}$ is $R$-linear, $w^{0}$ is a weight structure on ${\underline{C}^{0}}$, and $t$ is a $t$-structure on $\underline{C}$ strictly orthogonal to $w^{0}$ then for any weak Serre subcategory $\underline{A}$ of $R-\operatorname{Mod}$ one can specify a subcategory $\underline{C}_{\underline{A}}$ such that $t$ restricts to it.
@@ 2.1 Virtual $t$-truncations of (co)homological functors
(resp. $M\mapsto\operatorname{\operatorname{Im}}(H(w_{\geq n}M)\to H(w_{\geq n-1}M))$); here we take arbitrary choices of the corresponding weight truncations of $M$ and connect them using Proposition 1.2.4(4) in the case $g=\operatorname{id}_{M}$.
2. Let $H^{\prime}:\underline{C}\to\underline{A}$ be a homological functor. Then we will write $\tau_{\leq n}(H^{\prime})$ for the correspondence $M\mapsto\operatorname{\operatorname{Im}}(H^{\prime}(w_{\leq n}M)\to H^{\prime}(w_{\leq n+1}M))$ and $\tau_{\geq n}(H^{\prime})=M\mapsto\operatorname{\operatorname{Im}}(H^{\prime}(w_{\geq n-1}M)\to H^{\prime}(w_{\geq n}M))$ (here we take the same connecting arrows between weight truncations of $M$ as above).
3. Assume that $\underline{C}$ is a triangulated subcategory of a triangulated category $\underline{D}$. Then for any $M\in\operatorname{Obj}\underline{D}$ we will write $H_{M}=H_{M}^{\underline{C}}$ (resp. $H^{M}=H^{M}_{\underline{C}}$) for the restriction of the functor (co)represented by $M$ to $\underline{C}$ (thus $H_{M}$ and $H^{M}$ are functors from $\underline{C}$ into $\operatorname{Ab}$). Moreover, sometimes we will say that these functors are $\underline{D}$-Yoneda ones, and that $H_{M}$ (resp. $H^{M}$) is $\underline{D}$-(co)represented by $M$.
@@
    There exist natural transformations that yield a long exact sequence
@@
    (i.e., the result of applying this sequence to any object of $\underline{C}$ is a long exact sequence); the shift of this exact sequence by $3$ positions is given by composing the functors with $-[1]$.
@@
    Assume that there exists a $t$-structure $t$ that is left adjacent to $w$ (for certain $\underline{C}^{\prime}$ and $\underline{D}$ as in Definition 1.3.1). Then for any $M\in\operatorname{Obj}\underline{C}^{\prime}$ the functors $\tau_{\geq n}(H_{M})$ and $\tau_{\leq n}(H_{M})$ (where $H_{M}$ is defined in Definition 2.1.1(3)) are $\underline{D}$-represented (on $\underline{C}$) by $t_{\geq n}M$ and $t_{\leq n}M$ (see Proposition 1.1.5(3)), respectively.
@@
    Let $H^{op}$ be the cohomological functor obtained from $H$ by means of inverting arrows both in $\underline{C}$ and $\underline{A}$. Then there exist canonical isomorphisms $\tau^{op}_{\leq n}(H^{op})\cong(\tau_{\geq-n}(H))^{op}$ and $\tau^{op}_{\leq n}(H^{op})\cong(\tau_{\leq-n}(H))^{op}$; here $\tau^{op}$ denotes virtual $t$-truncations with respect to $w^{op}$.
@@
    Moreover, if there exists a $t$-structure $t$ on a triangulated category $\underline{C}^{\prime}$ that is orthogonal to $w$ (with respect to a triangulated category $\underline{D}$ containing $\underline{C}$ and $\underline{C}^{\prime}$), $\underline{A}=\operatorname{Ab}$, and the functor $H$ is $\underline{D}$-corepresented by an object $N$ of $\underline{C}^{\prime}$, then $\tau_{\geq n}(H^{\prime})$ is $\underline{D}$-corepresented by $t_{\geq n}N$ and $\tau_{\leq n}(H^{\prime})$ is $\underline{D}$-corepresented by $t_{\leq n}N$.
@@
Assertions 13 are essentially given by Theorem 2.3.1 of [Bon10b]; yet pay attention to Remark 1.2.3(3). Assertion 4 is given by Proposition 2.5.4(1) of ibid.; here one should apply Lemma 1.1.2 and recall Remark 1.3.4(1). Assertion 6 is easily seen to be dual to assertions 13.
Assertion 5 can be obtained by means of applying Proposition 1.2.4(1) quite straightforwardly. Firstly, reversing arrows converts any $m$-weight decomposition of $M\in\operatorname{Obj}\underline{C}$ into its $-m-1$-weight decomposition in $\underline{C}^{op}$ (with respect to $w^{op}$. Moreover, if we reverse arrows in (1.2.3) then we obtain a similar diagram in $\underline{C}^{op}$; if $m<n$ in the initial diagram then we obtain the unique morphism of triangles which corresponds to $g$ and connects (the dual) $-n-1$-weight decomposition of $M^{\prime}$ with the corresponding $-m-1$-weight decomposition of $M$. Thus if we reverse in the definitions $\tau_{\geq-n}(H)$ and $\tau_{\leq-n}(H)$ (for any $n\in{\mathbb{Z}}$) then we obtain the morphisms that compute $\tau^{op}_{\leq n}(H^{op})$ and $\tau^{op}_{\geq n}(H^{op})$, respectively. ∎
@@
Now let us list a few easy properties of $R$-linear categories, where $R$ is a commutative ring.
@@
Assertions 1 is very easy; this is Lemma 3.2.8(1) of [Bon24]. Assertion 2 follows from it immediately.
@@ 2.2 On weight range
    The (representable) functor $H_{M}=\underline{C}(-,M):\underline{C}^{op}\to\operatorname{Ab}$ if of weight range $\geq m$ if and only if $M\in\underline{C}_{w\geq m}$.
@@
Assertion 9 immediately follows from Lemma 2.1.4 of [Bon21].
@@
1. Roughly, the statements above say that virtual $t$-truncations of functors behave as if they corresponded to $t$-truncations of objects in a certain triangulated "category of functors" (whence the name; certainly, another justification of this idea is provided by the existence of orthogonal $t$-structures statements that will be proved below).
2. In particular, one can "slice" any functor of weight range $[m,n]$ for $m\leq n$ into "pieces" of weight $[i,i]$ for $m\leq i\leq n$. Next, composing a "slice" of weight range $[i,i]$ with $[i]$ one obtains a pure functor.
@@
1. If $H^{\prime}$ is of weight range $\geq n$, $H^{\prime\prime}$ is of weight range $\leq n-1$, then there exists a canonical isomorphism of chains of morphisms of functors (2.2.2)$\cong\tau_{\geq n}(H)\stackrel{{\scriptstyle a_{n}}}{{\to}}H\stackrel{{\scriptstyle b_{n-1}}}{{\to}}\tau_{\leq n-1}(H)$; the latter one is a part of the sequence (2.1.1).
@@
III. Assume that $\Phi:\underline{C}^{op}\times\underline{C}^{\prime}\to\underline{A}$ is a duality (in the sense of Definition 1.1.1) and $H=\Phi(-,M)$ for some object $M$ of $\underline{C}^{\prime}$. Then the transformation $a_{n}$ corresponding to $H$ equals $\Phi(-,f)$ for some $\underline{C}^{\prime}$-morphism $f:L\to M$ if and only if $b_{n-1}=\Phi(-,g)$ for a $\underline{C}^{\prime}$-morphism $g:M\to R$. Moreover, the morphisms $L\stackrel{{\scriptstyle f}}{{\to}}M\stackrel{{\scriptstyle g}}{{\to}}R$ can be completed to a distinguished triangle in $\underline{C}^{\prime}$.
@@
Then the sequence $\Phi(-,L\stackrel{{\scriptstyle f}}{{\to}}M\stackrel{{\scriptstyle g}}{{\to}}R)$ clearly satisfies our assumptions on (2.2.2) in assertion II. Then assertion II.2 implies that $b_{n-1}=\Phi(-,g)$. We also obtain the "moreover" statement in our assertion.
Conversely, if $b_{n-1}=\Phi(-,g)$ then we also can complete $g$ to a distinguished triangle of the form (2.2.4). We apply assertion II.2 once again to obtain $a_{n}=\Phi(-,f)$. ∎
@@
2.3 Some properties and restrictions of strictly orthogonal $t$-structures

### v2

@@ 1.2 Some basics on weight structures
Assume that $\mathcal{A}$ and $\mathcal{B}$ are strictly full triangulated subcategories of $\underline{C}$.
Then the couple $D=(\mathcal{A},\mathcal{B})$ is a semi-orthogonal decomposition of $\underline{C}$ (or just gives a decomposition of $\underline{C}$) if $\operatorname{Obj}\mathcal{B}\perp\operatorname{Obj}\mathcal{A}$ and $\operatorname{Obj}\mathcal{B}\star\operatorname{Obj}\mathcal{A}=\operatorname{Obj}\underline{C}$. [^3] Recall that $\operatorname{Obj}\mathcal{B}\star\operatorname{Obj}\mathcal{A}$ is the class of all extensions of objects of $\mathcal{B}$ by objects of $\mathcal{A}$.
@@
    The full subcategories of $\underline{C}$ corresponding to $C_{2}$ and $C_{1}$ give a semi-orthogonal decomposition of $\underline{C}$.
@@
    $(C_{2},C_{1})$ is a weight structure $w$ and ${\underline{Hw}}=\{0\}$.
@@
    $(C_{2},C_{1})$ is a weight structure $w$ and the triangle (1.2.1) is functorially determined by $M$.
@@
    $(C_{2},C_{1})$ is a weight structure $w$ and the triangle (1.2.1) is determined by $M$ up to a non-canonical isomorphism.
@@
The equivalence of conditions 3, 2, and 4 can be easily obtained by means of combining Propositions 3.4(4) and 3.2(1,2) of [BoV19] (pay attention to Remark 1.1.4(2) above!).
The equivalence of conditions 1 and 3 is very well-known; see Proposition 2.5 of [Bon24b].
@@
Lastly, assume that condition 5 is fulfilled. Then any weight decomposition triangle is clearly a $1$-weight decomposition triangle as well. Consequently, the uniqueness provided by Proposition 1.2.4(4) yields condition 7; cf. Remark 1.2.6 of [Bon24a] where compositions of the corresponding morphisms between decomposition triangles is discussed. ∎
@@
1. Roughly, semi-orthogonal decompositions are (essentially) the weight structures that are also $t$-structures. One may also characterize semi-orthogonal decompositions both as shift-stable weight structures and as shift-stable $t$-structures.
2. This notion is obviously self-dual (as well). That is, permuting the classes in the couple yields a semi-orthogonal decomposition in the category $\underline{C}^{op}$; cf. Remark 1.1.4(1) and Proposition 1.2.4(1).
3. Below we will also need the following simple criterion: if $\underline{C}$ is smashing (cf. Theorem 0.1) then a semi-orthogonal decomposition $w$ is smashing if and only if the right adjoint functor provided by condition 3 of Proposition 1.2.6 respects coproducts; see Proposition 3.4(5) of [BoV19].
@@ 1.3 On orthogonal and adjacent structures
1. We will say that $w$ and $t$ are orthogonal (in $\underline{D}$) if $\underline{C}_{w\leq 0}\perp_{\underline{D}}\underline{C}^{\prime}_{t\geq 1}$ and $\underline{C}_{w\geq 0}\perp_{\underline{D}}\underline{C}^{\prime}_{t\leq-1}$.
@@
Let us now relate the notion of adjacent structures to the definition given in [Bon10a].
@@
Proposition 1.3.3 says that our definition of adjacent structures is essentially equivalent to the original Definition 4.4.1 of [Bon10a].
@@
Assume that $w$ and $t$ satisfy the assumptions of Definition 1.3.1, $t$ is strictly orthogonal to $w$ (note that this is condition follows from orthogonality whenever $\underline{C}^{\prime}=\underline{C}$; see Theorem 2.3.3(I) below), and the couple $(\underline{C}_{w\geq 0},\underline{C}_{w\leq 0})$ is a semi-orthogonal decomposition (see Definition 1.2.5). Then $(\underline{C}^{\prime}_{t\leq 0},\underline{C}^{\prime}_{t\geq 0})$ is a semi-orthogonal decomposition as well.
@@
Applying Proposition 1.2.6 (see conditions 4 and 1 in it) along with the definition of strict orthogonality we obtain that $\underline{C}^{\prime}_{t\geq 0}=\underline{C}^{\prime}_{t\geq 0}[1]$. Hence $(\underline{C}^{\prime}_{t\leq 0},\underline{C}^{\prime}_{t\geq 0})$ is a semi-orthogonal decomposition indeed; see condition 2 in Proposition 1.2.6. ∎
@@ 2 On virtual $t$-truncations and their relation to orthogonal $t$-structures
In §2.3 we prove that if $\underline{C}^{\prime}\subset\underline{C}$ then for a weight structure $w$ on $\underline{C}$ there exists at most one $t$-structure on $\underline{C}^{\prime}$ (strictly) orthogonal to it; see Theorem 2.3.3. Moreover, if ${\underline{C}^{0}}$ is $R$-linear, $w^{0}$ is a weight structure on ${\underline{C}^{0}}$, and $t$ is a $t$-structure on $\underline{C}$ strictly orthogonal to $w^{0}$ then for any weak Serre subcategory $\underline{A}$ of $R-\operatorname{Mod}$ one can specify a subcategory $\underline{C}_{\underline{A}}$ such that $t$ restricts to it.
@@ 2.1 Virtual $t$-truncations of (co)homological functors
Now we will give some definitions central for this paper.
@@
(resp. $M\mapsto\operatorname{\operatorname{Im}}(H(w_{\geq n}M)\to H(w_{\geq n-1}M))$). Here we take arbitrary choices of the corresponding weight truncations of $M$ and connect them by means of applying Proposition 1.2.4(4) in the case $g=\operatorname{id}_{M}$.
2. Dually, let $H^{\prime}:\underline{C}\to\underline{A}$ be a homological functor. Then we will write $\tau_{\leq n}(H^{\prime})$ for the correspondence $M\mapsto\operatorname{\operatorname{Im}}(H^{\prime}(w_{\leq n}M)\to H^{\prime}(w_{\leq n+1}M))$ and $\tau_{\geq n}(H^{\prime})=M\mapsto\operatorname{\operatorname{Im}}(H^{\prime}(w_{\geq n-1}M)\to H^{\prime}(w_{\geq n}M))$ (here we take the same connecting arrows between weight truncations of $M$ as above).
3. Assume that $\underline{C}$ is a triangulated subcategory of a triangulated category $\underline{D}$. Then for any $M\in\operatorname{Obj}\underline{D}$ we will write $H_{M}=H_{M}^{\underline{C}}$ (resp. $H^{M}=H^{M}_{\underline{C}}$) for the restriction of the functor (co)represented by $M$ to $\underline{C}$ (thus $H_{M}$ and $H^{M}$ are functors from $\underline{C}$ into $\operatorname{Ab}$). Moreover, sometimes we will say that $H_{M}$ (resp. $H^{M}$) is $\underline{D}$-(co)represented by $M$.
Furthermore, for a $\underline{D}$-morphism $f:M_{1}\to M_{2}$ we will write $H_{f}^{\underline{C}}$ for the transformation $H_{M_{1}}^{\underline{C}}\to H_{M_{2}}^{\underline{C}}$ corresponding to $f$.
@@
    There exists a long exact sequence
@@
    of functors (i.e., the result of applying this sequence to any object of $\underline{C}$ is a long exact sequence). The shift of this exact sequence by $3$ positions to the right is given by composing the functors with $[1]$.
@@
    Assume that there exists a $t$-structure $t$ on $\underline{C}^{\prime}$ that is orthogonal to $w$ (for certain $\underline{C}^{\prime}$ and $\underline{D}$ as in Definition 1.3.1). Then for any $M\in\operatorname{Obj}\underline{C}^{\prime}$ the functors $\tau_{\geq n}(H_{M})$ and $\tau_{\leq n}(H_{M})$ are $\underline{D}$-represented (on $\underline{C}$) by $t_{\geq n}M$ and $t_{\leq n}M$ (see Definition 2.1.1(3) and Proposition 1.1.3(3)), respectively.
@@
    Let $H^{op}$ be the cohomological functor obtained from $H$ by means of inverting arrows both in $\underline{C}$ and $\underline{A}$. Then there exist canonical isomorphisms $\tau^{op}_{\leq n}(H^{op})\cong(\tau_{\geq-n}(H))^{op}$ and $\tau^{op}_{\geq n}(H^{op})\cong(\tau_{\leq-n}(H))^{op}$; here $\tau^{op}$ denotes virtual $t$-truncations with respect to $w^{op}$.
@@
    Moreover, if there exists a $t$-structure $t$ on a triangulated category $\underline{C}^{\prime}$ that is orthogonal to $w$ (in a triangulated category $\underline{D}$ containing $\underline{C}$ and $\underline{C}^{\prime}$), $\underline{A}=\operatorname{Ab}$, and the functor $H$ is $\underline{D}$-corepresented by an object $N$ of $\underline{C}^{\prime}$, then $\tau_{\geq n}(H^{\prime})$ is $\underline{D}$-corepresented by $t_{\geq n}N$ and $\tau_{\leq n}(H^{\prime})$ is $\underline{D}$-corepresented by $t_{\leq n}N$.
@@
Assertions 13 are essentially given by Theorem 2.3.1 of [Bon10b]; yet pay attention to Remark 1.2.3(3). Assertion 4 is given by Proposition 2.5.4(1) of ibid. (see Remark 6.2 below for some detail). Assertion 6 is easily seen to be dual to assertions 13.
Assertion 5 can be obtained by means of applying Proposition 1.2.4(1) quite straightforwardly. Firstly, reversing arrows converts any $m$-weight decomposition of $M\in\operatorname{Obj}\underline{C}$ into its $-m-1$-weight decomposition in $\underline{C}^{op}$ (with respect to $w^{op}$). Moreover, if we reverse arrows in (1.2.3) then we obtain a similar diagram in $\underline{C}^{op}$; if $m<n$ in the initial diagram then we obtain the unique morphism of triangles which corresponds to $g$ and connects (the dual) $-n-1$-weight decomposition of $M^{\prime}$ with the corresponding $-m-1$-weight decomposition of $M$. Thus if we reverse in the definitions $\tau_{\geq-n}(H)$ and $\tau_{\leq-n}(H)$ (for any $n\in{\mathbb{Z}}$) then we obtain the morphisms that compute $\tau^{op}_{\leq n}(H^{op})$ and $\tau^{op}_{\geq n}(H^{op})$, respectively. ∎
@@
Now let us list a few easy properties of $R$-linear categories.
Throughout this paper $R$ will be an associative commutative unital ring.
@@
Assertions 1 is very easy; this is Lemma 3.11(1) of [Bon24b]. Assertion 2 follows from it immediately.
@@ 2.2 On weight range
    The (representable) functor $H_{M}=\underline{C}(-,M):\underline{C}^{op}\to\operatorname{Ab}$ is of weight range $\geq m$ if and only if $M\in\underline{C}_{w\geq m}$.
@@
Assertion 9 easily follows assertion 5; see also Lemma 2.1.4 of [Bon21].
@@
1. Roughly, the statements above say that virtual $t$-truncations of functors behave as if they corresponded to $t$-truncations of objects in a certain triangulated ’category of functors’ (whence the name; certainly, another justification of this idea is provided by the existence of orthogonal $t$-structures statements that will be proved below).
2. In particular, one can ’slice’ any functor of weight range $[m,n]$ for $m\leq n$ into ’pieces’ of weight $[i,i]$ for $m\leq i\leq n$. Next, composing a ’slice’ of weight range $[i,i]$ with $[i]$ one obtains a pure functor.
@@
1. If $H^{\prime}$ is of weight range $\geq n$, $H^{\prime\prime}$ is of weight range $\leq n-1$, then there exists a canonical isomorphism of chains of morphisms of functors
(2.2.2)$\cong\tau_{\geq n}(H)\stackrel{{\scriptstyle a_{n}}}{{\to}}H\stackrel{{\scriptstyle b_{n-1}}}{{\to}}\tau_{\leq n-1}(H)$;
here the latter chain is a part of the sequence (2.1.1).
@@
III. Assume that $\underline{C},\underline{C}^{\prime}\subset\underline{D}$ and $H=H_{M}^{\underline{C}}$ for some object $M$ of $\underline{C}^{\prime}$. Then the transformation $a_{n}$ corresponding to $H$ can be presented as $H^{\underline{C}}_{f}$ (that is, the the restriction to $\underline{C}$ of $\underline{D}(-,f)$; see Definition 2.1.1(3)) for some $\underline{C}^{\prime}$-morphism $f:L\to M$ if and only if $b_{n-1}$ is of the form $H^{\underline{C}}_{g}$ for some $\underline{C}^{\prime}$-morphism $g:M\to R$.
Moreover, the morphisms $L\stackrel{{\scriptstyle f}}{{\to}}M\stackrel{{\scriptstyle g}}{{\to}}R$ can be completed to a distinguished triangle in $\underline{C}^{\prime}$.
@@
Then the restriction of the transformation sequence $\underline{D}(-,L\stackrel{{\scriptstyle f}}{{\to}}M\stackrel{{\scriptstyle g}}{{\to}}R)$ to $\underline{C}$ clearly satisfies our assumptions on (2.2.2) in assertion II. Then assertion II.2 implies that $b_{n-1}$ is the corresponding restriction of $\underline{D}(-,g)$. We also obtain the "moreover" statement in our assertion.
Conversely, if $b_{n-1}\cong H^{\underline{C}}_{g}$ then we also can complete $g$ to a distinguished triangle of the form (2.2.4). We apply assertion II.2 once again to obtain that $a_{n}\cong H^{\underline{C}}_{f}$. ∎
@@
2.3 Strictly orthogonal $t$-structures (and their restrictions)